OpenStudy (anonymous):

could someone help me with Domain Values of Composite Functions (alg. 2.2)

OpenStudy (anonymous):

What is your question?

OpenStudy (anonymous):

Let f(x)=1/x and g(x)=x^2 -2x , what two numbers are NOT in the domain of F of G

OpenStudy (anonymous):

What are your choices?

OpenStudy (anonymous):

i dont have any :(

OpenStudy (anonymous):

ah ohtay one sec k?

OpenStudy (anonymous):

lol okay

OpenStudy (anonymous):

Are you doing Parabola's & Directrix's?

OpenStudy (anonymous):

i have no idea im honestly confused with my math right now

OpenStudy (anonymous):

Are you in FLVS?

OpenStudy (anonymous):

no lol idk what that is but im in treca

OpenStudy (anonymous):

ah... :3 one moment, let me do some research

OpenStudy (anonymous):

okay :3

OpenStudy (anonymous):

otay, Do you understand that the domain is every number that x is allowed to be?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

That, the only number it can't be is 0, because it real numbers, you can't divide be 0?

OpenStudy (anonymous):

okaay...

OpenStudy (anonymous):

So, becuase of this fact, your first number that is restricted by the domain, is 0

OpenStudy (anonymous):

so one of the numbers is 0?

OpenStudy (anonymous):

yeppers cx

OpenStudy (anonymous):

and 2 would be the other?

OpenStudy (anonymous):

Nevermind, it's 2 lol

OpenStudy (anonymous):

@lexikaylynn

OpenStudy (anonymous):

lol i thought so

OpenStudy (anonymous):

my bad cx... lol I love ADTR just sayin cx

OpenStudy (anonymous):

Love them lol <3

OpenStudy (anonymous):

& asking alex, & Black veil Brides xD

OpenStudy (anonymous):

yep big fan of all lol and om&m and Attack Attack!

OpenStudy (anonymous):

Attack Attack! </3 they broke up xc

OpenStudy (anonymous):

okay i need help on another question but its the same concept as the last and whaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaa nooooooooooooooooooooooooo!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

Post it, & yep, in 2012

OpenStudy (anonymous):

why!!!!!!!!!!!! i loved em ha and okay hold on

OpenStudy (anonymous):

They wanted to settle down. LoL that's why Austin sings for OM&M now cx

OpenStudy (anonymous):

i knew that lol well i forgot that Austin was in aa lol FAIL!!!!!

OpenStudy (anonymous):

He's not in AA, He's in OM&M.

OpenStudy (anonymous):

i know that lol

OpenStudy (anonymous):

cx <3

OpenStudy (anonymous):

\[given f(x)=\sqrt{9x} and g(x)=1/x-1 which value is \in the domain of f of g\]

OpenStudy (anonymous):

well that didnt work loke plained

OpenStudy (anonymous):

given \[f(x)=\sqrt{9x}\] and \[g(x)=1/x-1\] which value is in the domain of (f of g)

OpenStudy (anonymous):

A) -5 B) 1 C)-6 D)5

OpenStudy (anonymous):

A

OpenStudy (anonymous):

nope it was wrong

OpenStudy (anonymous):

Go to someone else... I suck

OpenStudy (anonymous):

ha its okay lol

OpenStudy (anonymous):

meh

OpenStudy (anonymous):

the only other thing would be graphing piecewise functions

OpenStudy (anonymous):

I have no idea... sorry

OpenStudy (anonymous):

lol its okay

OpenStudy (anonymous):

fan meh back or somethin

OpenStudy (anonymous):

i did